In a crank and slotted lever mechanism, the driving crank is 35 mm long, and the time ratio of cutting stroke to return stroke is 1.6. If the length of working stroke of the ram is 110 mm, find the distance between the fixed centers, and the slotted lever length.


1
Expert's answer
2021-05-28T02:26:18-0400

"cos\frac{\alpha}{2}=\frac{35}{110} \implies \frac{\alpha}{2}=cos^{-1}\frac{35}{110}=71.45^0"

"L=2 AR sin(\frac{\alpha}{2})"

"110=2 AR sin(71.45)"

"AR=\frac{110}{2 * sin 71.45}=58.01 mm"